The 1,000 accounts receivable of Baker Company have a total book value of $30,000 (Average book value = $30). Wendy Duffo, CPA, has selected and audited a sample of 50 accounts with the following mean values:

1. Book value of $29.00
2. Audited value of $29.50.

TB MC Qu. 09-20 What is the estimated total audited value...

What is the estimated total audited value using mean-per-unit sampling?

Multiple Choice

  • $29,000.

  • $29,500.

  • $30,500.

  • $30,517.

  • The 4,000 accounts receivable of Miller Company have a total book value of $150,000. A CPA has selected and audited a sample of 100 accounts with a total book value of $3,700 and an audited value of $3,800.

    TB MC Qu. 09-13 Using the mean-per-unit estimation...

    Using the mean-per-unit estimation technique, the estimated total audited value of the population is:

    Multiple Choice

  • $150,000.

  • $152,000.

  • $154,000.

  • $154,054.
Answer #1

Answer:

Total Audited value:

as per the difference estimation technique

audit value = book value + error.

book value = $ 30,000

error = (sample audit value - sample book value ) / sample size) x total population

= (29.50 - 29) / 50) x 1000 accounts

= $ 500

audit value of population = $ 30000 + $ 500 = $ 30500

---------------------

Total audited value for 100 accounts = 3800

Mean vale = 3800/100 = 38

Total value per mean per unit = 38* 4000

= $ 152,000

----------------------

If you have any query regarding the answer please as me in the comment I am here for help you. Please do not direct thumbs down just ask if you have any query. And if you like my work then please appreciates with up vote
Thank you Very much.